When you drink cold water, your body must expend metabolic energy in order to maintain normal body temperature (37.0°C) by warming up the water in your stomach. Could drinking ice water, then, substitute for exercise as a way to "burn calories?" Suppose you expend 490. kilocalories during a brisk hour-long walk. How many liters of ice water (0°C) would you have to drink in order to use up 490. kilocalories of metabolic energy? For comparison, the stomach can hold about 1 liter.
Answer #1

given
Q = 490 kcal

= 490*10^3*4.186 J

we know, specific heat of water, C = 4186 J/(kg C)

let m is the mass of ice water required.

use, Q = m*C*delta_t

==> m = Q/(C*delta_T)

= 490*10^3*4.186/(4186*(37 - 0) )

= 13.2 kg

volume of ice water need to have = 13.2 L <<<<<<<<<---------------Answer

(since 1 L of water has mass 1 kg)